ADNAN MUHAMMED

Grade 12,

Iodine Molecue dissociates into atoms after absorbing light of 4500 A. If oone quantum of radiation is absorbed by each molecule, calculate K.E. of iodine atoms.

Iodine Molecue dissociates into atoms after absorbing light of 4500 A. If oone quantum of radiation is absorbed by each molecule, calculate K.E. of iodine atoms.

Grade:11

1 Answers

deeksha sharma
13 years ago

WAVELTH=4500*10-10 m

energy=hc/wavelength=6.62*10-34 *3*108 /4500*10-10 =4.4*10-19 J

half of this energy goes to each atom i.e. 2.2*10-19 J

bond energy of iodine= 240kj/mol = 3.9*10-19 J/atom

for one atom=1.95*10-19 J

ke of iodine atoms= (2.2-1.95)*10-19 J =0.25*10-19 J=2.5*10-20 J